OpenStudy (anonymous):

Find the sum s8 for the geometric series 6+0.6+0.06+0.006+... A) 6.6666666 B) 0.1 C) -0.00000001 D) 3

OpenStudy (tkhunny):

The very first piece, "6", makes all but one answer pretty silly.

OpenStudy (anonymous):

Do you see any patterns? Geometric sum just means add all those numbers up!

OpenStudy (anonymous):

So its A?

OpenStudy (anonymous):

I figured it would be A because of the numbers, I just wasn't sure how to find it!

OpenStudy (anonymous):

6 + 0.6 = 6.6 6.6 + 0.06 = 6.66 6.66 + 0.006 = 6.666 and so on....